OpenStudy (anonymous):

Find S5 for each series

OpenStudy (anonymous):

\[S_{5}\]

OpenStudy (anonymous):

\[1 + \frac{ 1 }{ 2} + \frac{ 1 }{ 4 } + \frac{ 1 }{ 8 } + ...\]

OpenStudy (anonymous):

@ikram002p @mathstudent55 @SolomonZelman

OpenStudy (ikram002p):

S_5 means sumation ?

OpenStudy (ikram002p):

or the sumation untill the term 5 ?

OpenStudy (anonymous):

yes.

OpenStudy (ikram002p):

ok lol sum them :D did u give it a try ?

OpenStudy (anonymous):

yeah I get 1.875

OpenStudy (ikram002p):

1+1/2+1/4+1/8=1.875 1+1/2+1/4+1/8+1/16 = 31/16= 1.9375

OpenStudy (anonymous):

Ohhh get it! the pattern is the denominator 2n! thanks!

OpenStudy (ikram002p):

no no not 2n its 2^n

OpenStudy (anonymous):

ahh I see, thanks for the help!

OpenStudy (ikram002p):

np , ur wlc :D